A new high school math project is being tested in the state. In the first year​ (year 1), there are 671 students involved. By the end of the third year​ (year 3), 12487 students are expected to be involved. Assuming the increase is​ linear, write the equation of the​ line, in​ slope-intercept form, representing the number of students participating in this project for each year x. Use this result to predict the number of students involved at the end of 5 years.
​Hint: Use the ordered pairs​ (1, 671​) and​ (3, 12487​).

Answer by Theo(13342)

the slope intercept form of a straight line equation is:
y = mx + b
m is the slope
b is the y-intercept.
you have two points to use.
they are:
(1,671)
(3,12487)
m is the slope and is equal to (12487 - 671) / (3 - 1) = 5908.
the slope intercept form of the equation becomes:
y = 5908 * x + b
in this equation, x represents the year.
the first year becomes x = 0
the third year becomes x = 2
your coordinate points becomes (x,y) = (0,671) and (2,12487).
this a translation from (year, number of students) = (1,671) and (3,12487)
when year = 1, x = 0
when year = 3, x = 2
using the point (x,y) = (0,671), we find b to be equal to 671, because the y intercept is the value of y when x is equal to 0.
if we had used the point (2,12487), we would have solved for b as follows:
y = 5908 * x + b becomes 12487 = 5908 * 2 + b
we would solve for b to get b = 12487 - 2 * 5908 = 12487 - 11816 = 671.
both points lead to the same value of b, as they should.
the equation becomes y = 5908 * x + 671.
when x = 0, y = 671
when x = 1, y = 5908 * 1 + 671 = 6579
when x = 2, y = 5908 * 2 + 671 = 12487.
this equation can be graphed as shown below:

$$$